OpenStudy (anonymous):

Find the sum of the first 9 terms of the sequence. Show all work for full credit. 2, -7, -16, -25

OpenStudy (anonymous):

sorry, x-9. Keep subtracting 9 until you get 9 terms

OpenStudy (anonymous):

Then add up the terms

OpenStudy (anonymous):

2, -7, -16, -25, -34, -43, -52, -61, -70

OpenStudy (anonymous):

great, now add them up

OpenStudy (anonymous):

-306?

OpenStudy (anonymous):

@10115658

OpenStudy (anonymous):

yep, good job

OpenStudy (anonymous):

Thats it ?

OpenStudy (anonymous):

@10115658

OpenStudy (anonymous):

ya that should be the answer, quick and easy

OpenStudy (anonymous):

wanna help on another?

OpenStudy (anonymous):

Start a new question and i'll have a look

OpenStudy (anonymous):

Find the sum of the geometric sequence. three divided by two, three divided by eight, three divided by thirty two, three divided by one hundred and twenty eight, three divided by five hundred and twelve one divided by two hundred and fifty six one divided by sixteen one thousand twenty three divided by five hundred and twelve 341

OpenStudy (triciaal):

first term a = 2 difference between terms, d = -9 looking for 9th term so use n = 9 formula sum of terms of an arithmetic progression sum = a(n) = a(1) + a(n - 1) d a(9) = 2 + (9-1)(-9) 2 - 72 = -70